# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ IV.2.1 $\overline{B}\to D\overline{T}$ decay amplitudes
$\Gamma(\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s})=\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}}).$ (45)
The rates in Eqs. (44) and (45) are expected to be larger than those in Eq. (43) in the same $\Delta S$ transition. as the last two sets have contributions from external $W$ amplitudes, $T_{\bar{T}}$ or $T^{\prime}_{\bar{T}}$.
Note that in ref. Qin:2022nof the amplitudes of the following four modes in $\Delta S=-1$ transitions in scenario II, namely $B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}s}$, $\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}s}$, $B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}ss}$, $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}ss}$, were also given. Their results agree with those shown in Table 8, but with $C_{\bar{T}}$ replaced with $CV_{cb}V^{*}_{cs}$, while the rest in Table 8 and all in Table 7 are new.
The final state $\overline{T}$ can further decay to $\bar{D}P$ and $\bar{D}S$, or they can connect to these states virtually. One can obtain their full amplitudes using results in the section and those in Sec. III.
@@ IV.2.2 $\overline{B}\to\overline{D}T$ decay amplitudes
# Mode $A(\overline{B}_{q}\to\overline{D}T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]})$ Mode $A(\overline{B}_{q}\to\overline{D}T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]})$
$4^{*}$ $B^{-}\to D^{-}T^{0}_{cs}$ $C_{\bar{D}T}$ $4^{*}$ $\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{cs}$ $-C_{\bar{D}T}$
5 $\overline{B}{}^{0}\to D_{s}^{-}T^{+}_{c}$ $\frac{C_{\bar{D}T}}{\sqrt{2}}$ 5 $\overline{B}{}_{s}^{0}\to D^{-}T^{+}_{c}$ $-\frac{1}{\sqrt{2}}C_{\bar{D}T}$
6 $B^{-}\to D_{s}^{-}T^{0}_{c}$ $-\frac{1}{\sqrt{2}}C_{\bar{D}T}$ 6 $\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{0}_{c}$ $\frac{1}{\sqrt{2}}C_{\bar{D}T}$
7 $\overline{B}{}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c}$ $-\frac{1}{\sqrt{2}}(2T_{\bar{D}}+C_{\bar{D}T})$ 7 $\overline{B}{}_{s}^{0}\to D^{-}T^{\prime\prime+}_{c}$ $\frac{1}{\sqrt{2}}(C_{\bar{D}T}+2E_{\bar{D}\bar{B}})$
8 $B^{-}\to D_{s}^{-}T^{\prime\prime 0}_{c}$ $-\frac{1}{\sqrt{2}}(2T_{\bar{D}}+C_{\bar{D}T})$ 8 $\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{\prime\prime 0}_{c}$ $\frac{1}{\sqrt{2}}(C_{\bar{D}T}+2E_{\bar{D}\bar{B}})$
9 $\overline{B}{}_{s}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c\bar{s}}$ $-\sqrt{2}(T_{\bar{D}}-E_{\bar{D}\bar{B}})$
# Mode $A^{\prime}(\overline{B}_{q}\to\overline{D}T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]})$ Mode $A^{\prime}(\overline{B}_{q}\to\overline{D}T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]})$
$2^{*}$ $\overline{B}{}^{0}\to D_{s}^{-}T^{+}_{c\bar{s}}$ $-\frac{1}{\sqrt{2}}C^{\prime}_{\bar{D}T}$ $2^{*}$ $\overline{B}{}_{s}^{0}\to D^{-}T^{+}_{c\bar{s}}$ $\frac{1}{\sqrt{2}}C^{\prime}_{\bar{D}T}$
$3^{*}$ $B^{-}\to D_{s}^{-}T^{0}_{c\bar{s}}$ $C^{\prime}_{\bar{D}T}$ $3^{*}$ $\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{0}_{c\bar{s}}$ $-C^{\prime}_{\bar{D}T}$
6 $B^{-}\to D^{-}T^{0}_{c}$ $\frac{1}{\sqrt{2}}C^{\prime}_{\bar{D}T}$ 6 $\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{c}$ $-\frac{1}{\sqrt{2}}C^{\prime}_{\bar{D}T}$
7 $\overline{B}{}^{0}\to D^{-}T^{\prime\prime+}_{c}$ $-\sqrt{2}(T^{\prime}_{\bar{D}}-E^{\prime}_{\bar{D}\bar{B}})$ 8 $B^{-}\to D^{-}T^{\prime\prime 0}_{c}$ $-\frac{1}{\sqrt{2}}(2T^{\prime}_{\bar{D}}+C^{\prime}_{\bar{D}T})$
8 $\overline{B}{}^{0}\to\overline{D}^{0}T^{\prime\prime 0}_{c}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{D}T}+2E^{\prime}_{\bar{D}\bar{B}})$
9 $\overline{B}{}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c\bar{s}}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{D}T}+2E^{\prime}_{\bar{D}\bar{B}})$ 9 $\overline{B}{}_{s}^{0}\to D^{-}T^{\prime\prime+}_{c\bar{s}}$ $-\frac{1}{\sqrt{2}}(2T^{\prime}_{\bar{D}}+C^{\prime}_{\bar{D}T})$
@@
# Mode $A(\overline{B}_{q}\to\overline{D}T_{cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}})$ Mode $A(\overline{B}_{q}\to\overline{D}T_{cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}})$
$4^{\prime*}$ $\overline{B}{}^{0}\to D^{-}T^{+}_{cs}$ $\sqrt{2}C_{\bar{D}T}$ $5^{\prime*}$ $B^{-}\to D^{-}T^{0}_{cs}$ $C_{\bar{D}T}$
$5^{\prime*}$ $\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{cs}$ $C_{\bar{D}T}$ $6^{\prime*}$ $B^{-}\to\overline{D}^{0}T^{-}_{cs}$ $\sqrt{2}C_{\bar{D}T}$
$13^{\prime}$ $\overline{B}{}^{0}\to D_{s}^{-}T^{\prime+}_{c}$ $-\frac{\sqrt{3}}{2}C_{\bar{D}T}$ $13^{\prime}$ $\overline{B}{}_{s}^{0}\to D^{-}T^{\prime+}_{c}$ $-\frac{\sqrt{3}}{2}C_{\bar{D}T}$
$14^{\prime}$ $B^{-}\to D_{s}^{-}T^{\prime 0}_{c}$ $-\frac{\sqrt{3}}{2}C_{\bar{D}T}$ $14^{\prime}$ $\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{\prime 0}_{c}$ $-\frac{\sqrt{3}}{2}C_{\bar{D}T}$
$15^{\prime}$ $\overline{B}{}_{s}^{0}\to D_{s}^{-}T^{\prime+}_{c\bar{s}}$ $-C_{\bar{D}T}$
$16^{\prime}$ $\overline{B}{}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c}$ $\frac{1}{2}(4T_{\bar{D}}+C_{\bar{D}T})$ $16^{\prime}$ $\overline{B}{}_{s}^{0}\to D^{-}T^{\prime\prime+}_{c}$ $\frac{1}{2}(C_{\bar{D}T}+4E_{\bar{D}\bar{B}})$
$17^{\prime}$ $B^{-}\to D_{s}^{-}T^{\prime\prime 0}_{c}$ $\frac{1}{2}(4T_{\bar{D}}+C_{\bar{D}T})$ $17^{\prime}$ $\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{\prime\prime 0}_{c}$ $\frac{1}{2}(C_{\bar{D}T}+4E_{\bar{D}\bar{B}})$
$18^{\prime}$ $\overline{B}{}_{s}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c\bar{s}}$ $2T_{\bar{D}}+C_{\bar{D}T}+2E_{\bar{D}\bar{B}}$
# Mode $A^{\prime}(\overline{B}_{q}\to\overline{D}T_{cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}})$ Mode $A^{\prime}(\overline{B}_{q}\to\overline{D}T_{cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}})$
$2^{\prime*}$ $\overline{B}{}^{0}\to D_{s}^{-}T^{+}_{c\bar{s}}$ $-\frac{C^{\prime}_{\bar{D}T}}{\sqrt{2}}$ $2^{\prime*}$ $\overline{B}{}_{s}^{0}\to D^{-}T^{+}_{c\bar{s}}$ $-\frac{C^{\prime}_{\bar{D}T}}{\sqrt{2}}$
$3^{\prime*}$ $B^{-}\to D_{s}^{-}T^{0}_{c\bar{s}}$ $C^{\prime}_{\bar{D}T}$ $3^{\prime*}$ $\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{0}_{c\bar{s}}$ $C^{\prime}_{\bar{D}T}$
$8^{\prime*}$ $\overline{B}{}_{s}^{0}\to D_{s}^{-}T^{+}_{c\bar{s}\bar{s}}$ $\sqrt{2}C^{\prime}_{\bar{D}T}$
$10^{\prime*}$ $\overline{B}{}^{0}\to D^{-}T^{+}_{c}$ $-\sqrt{\frac{2}{3}}C^{\prime}_{\bar{D}T}$ $11^{\prime*}$ $B^{-}\to D^{-}T^{0}_{c}$ $-\sqrt{\frac{2}{3}}C^{\prime}_{\bar{D}T}$
$11^{\prime*}$ $\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{c}$ $-\sqrt{\frac{2}{3}}C^{\prime}_{\bar{D}T}$ $12^{\prime*}$ $B^{-}\to\overline{D}^{0}T^{-}_{c}$ $\sqrt{2}C^{\prime}_{\bar{D}T}$
$13^{\prime}$ $\overline{B}{}^{0}\to D^{-}T^{\prime+}_{c}$ $\frac{1}{\sqrt{3}}C^{\prime}_{\bar{D}T}$ $14^{\prime}$ $B^{-}\to D^{-}T^{\prime 0}_{c}$ $\frac{1}{2\sqrt{3}}C^{\prime}_{\bar{D}T}$
$14^{\prime}$ $\overline{B}{}^{0}\to\overline{D}^{0}T^{\prime 0}_{c}$ $\frac{1}{2\sqrt{3}}C^{\prime}_{\bar{D}T}$
$15^{\prime}$ $\overline{B}{}^{0}\to D_{s}^{-}T^{\prime+}_{c\bar{s}}$ $\frac{1}{2}C^{\prime}_{\bar{D}T}$ $15^{\prime}$ $\overline{B}{}_{s}^{0}\to D^{-}T^{+}_{c\bar{s}}$ $\frac{1}{2}C^{\prime}_{\bar{D}T}$
$16^{\prime}$ $\overline{B}{}^{0}\to D^{-}T^{\prime\prime+}_{c}$ $2T^{\prime}_{\bar{D}}+C^{\prime}_{\bar{D}T}+2E_{\bar{D}\bar{B}}^{\prime}$ $17^{\prime}$ $B^{-}\to D^{-}T^{\prime\prime 0}_{c}$ $\frac{1}{2}(4T^{\prime}_{\bar{D}}+C^{\prime}_{\bar{D}T})$
$17^{\prime}$ $\overline{B}{}^{0}\to\overline{D}{}^{0}T^{\prime\prime 0}_{c}$ $\frac{1}{2}(C^{\prime}_{\bar{D}T}+4E^{\prime}_{\bar{D}\bar{B}})$
$18^{\prime}$ $\overline{B}{}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c\bar{s}}$ $\frac{1}{2}(C^{\prime}_{D.5}+4E^{\prime}_{\bar{D}\bar{B}})$ $18^{\prime}$ $\overline{B}{}_{s}^{0}\to D^{-}T^{\prime\prime+}_{c\bar{s}}$ $\frac{1}{2}(4T^{\prime}_{\bar{D}}+C^{\prime}_{\bar{D}T})$
@@
From Table 9 we see that the decay modes involving $T$ in ${\bf 6}$ in scenario I are governed by internal $W$ amplitudes $C_{\bar{D}T}$ or $C^{\prime}_{\bar{D}T}$. Consequently, the sizes of their decay rates are highly related. In particular, the modes involving flavor exotic states have the following relations on decay rates,
$\Gamma(B^{-}\to D^{-}T^{0}_{cs})$ $=$ $\Gamma(\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{cs})=\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(B^{-}\to D_{s}^{-}T^{0}_{c\bar{s}})$ (46)
$=$ $\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{0}_{c\bar{s}})=2\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to D_{s}^{-}T^{+}_{c\bar{s}})$
$=$ $2\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}_{s}^{0}\to D^{-}T^{+}_{c\bar{s}}).$

### v3

@@ IV.2.1 $\overline{B}\to D\overline{T}$ decay amplitudes
$\Gamma(\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s})\simeq\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}}),$ (66)
where the above approximations neglect penguin contributions with a 10% to 20% error, see Eq. (62). The rates in Eqs. (65) and (66) are expected to be larger than those in Eq. (64) in the same $\Delta S$ transition. as the last two sets have contributions from external $W$ amplitudes, $T_{\bar{T}}$ or $T^{\prime}_{\bar{T}}$.
Note that in ref. Qin:2022nof the amplitudes of the following four modes in $\Delta S=-1$ transitions in scenario II, namely $B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}s}$, $\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}s}$, $B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}ss}$, $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}ss}$, were also given. Their results agree with those shown in Table 8, but with $C_{\bar{T}}$ replaced with $CV_{cb}V^{*}_{cs}$ and neglecting $P_{\bar{T},1}$, as they do not consider penguin contributions. The rest in Table 8 and all in Table 7 are new. Note that ref. Qin:2022nof also used a different notation for the open-charmed tetraquark states, see Table 2 for a comparison. For example, in Table 2 of ref. Qin:2022nof , one finds
$A(B^{-}\to D^{+}T_{\bar{c}\bar{u}ds})=CV_{cb}V^{*}_{cs},$ (67)
which corresponds to
$A(B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}s})=C_{\bar{T}}+P_{\bar{T},1},$ (68)
in Table 8, but with $C_{\bar{T}}$ replaced with $CV_{cb}V^{*}_{cs}$ and neglecting $P_{\bar{T},1}$.
The final state $\overline{T}$ can further decay to $\bar{D}P$ and $\bar{D}S$, or they can connect to these states virtually. One can obtain their full amplitudes using the results in this section and those in Sec. III.
@@ IV.2.2 $\overline{B}\to\overline{D}T$ decay amplitudes
# Mode $A(\overline{B}_{q}\to\overline{D}T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]})$ # Mode $A(\overline{B}_{q}\to\overline{D}T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]})$
$4^{*}$ $B^{-}\to D^{-}T^{0}_{cs}$ $(C_{\bar{D}T}+P_{T})$ $4^{*}$ $\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{cs}$ $(C_{\bar{D}T}+P_{T})$
5 $\overline{B}{}^{0}\to D_{s}^{-}T^{+}_{c}$ $\frac{(C_{\bar{D}T}+P_{T})}{\sqrt{2}}$ 5 $\overline{B}{}_{s}^{0}\to D^{-}T^{+}_{c}$ $\frac{1}{\sqrt{2}}(C_{\bar{D}T}+P_{T})$
6 $B^{-}\to D_{s}^{-}T^{0}_{c}$ $\frac{1}{\sqrt{2}}(C_{\bar{D}T}+P_{T})$ 6 $\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{0}_{c}$ $\frac{1}{\sqrt{2}}(C_{\bar{D}T}+P_{T})$
7 $\overline{B}{}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c}$ $\frac{1}{\sqrt{2}}(2(T_{\bar{D}}+P_{\bar{D}})+(C_{\bar{D}T}+P_{T}))$ 7 $\overline{B}{}_{s}^{0}\to D^{-}T^{\prime\prime+}_{c}$ $\frac{1}{\sqrt{2}}((C_{\bar{D}T}+P_{T})+2(E_{\bar{D}\bar{B}}+PA_{\bar{B}}))$
8 $B^{-}\to D_{s}^{-}T^{\prime\prime 0}_{c}$ $\frac{1}{\sqrt{2}}(2(T_{\bar{D}}+P_{\bar{D}})+(C_{\bar{D}T}+P_{T}))$ 8 $\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{\prime\prime 0}_{c}$ $\frac{1}{\sqrt{2}}((C_{\bar{D}T}+P_{T})+2(E_{\bar{D}\bar{B}}+PA_{\bar{B}}))$
9 $\overline{B}{}_{s}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c\bar{s}}$ $\sqrt{2}((T_{\bar{D}}+P_{\bar{D}})-(E_{\bar{D}\bar{B}}+PA_{\bar{B}}))$
# Mode $A^{\prime}(\overline{B}_{q}\to\overline{D}T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]})$ # Mode $A^{\prime}(\overline{B}_{q}\to\overline{D}T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]})$
$2^{*}$ $\overline{B}{}^{0}\to D_{s}^{-}T^{+}_{c\bar{s}}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$ $2^{*}$ $\overline{B}{}_{s}^{0}\to D^{-}T^{+}_{c\bar{s}}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$
$3^{*}$ $B^{-}\to D_{s}^{-}T^{0}_{c\bar{s}}$ $(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$ $3^{*}$ $\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{0}_{c\bar{s}}$ $(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$
6 $B^{-}\to D^{-}T^{0}_{c}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$ 6 $\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{c}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$
7 $\overline{B}{}^{0}\to D^{-}T^{\prime\prime+}_{c}$ $\sqrt{2}((T^{\prime}_{\bar{D}}+P^{\prime}_{\bar{D}})-(E^{\prime}_{\bar{D}\bar{B}}+PA^{\prime}_{\bar{B}}))$ 8 $B^{-}\to D^{-}T^{\prime\prime 0}_{c}$ $\frac{1}{\sqrt{2}}(2(T^{\prime}_{\bar{D}}+P^{\prime}_{\bar{D}})+(C^{\prime}_{\bar{D}T}+P^{\prime}_{T}))$
8 $\overline{B}{}^{0}\to\overline{D}^{0}T^{\prime\prime 0}_{c}$ $\frac{1}{\sqrt{2}}((C^{\prime}_{\bar{D}T}+P^{\prime}_{T})+2(E^{\prime}_{\bar{D}\bar{B}}+PA^{\prime}_{\bar{B}}))$
9 $\overline{B}{}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c\bar{s}}$ $\frac{1}{\sqrt{2}}((C^{\prime}_{\bar{D}T}+P^{\prime}_{T})+2(E^{\prime}_{\bar{D}\bar{B}}+PA^{\prime}_{\bar{B}}))$ 9 $\overline{B}{}_{s}^{0}\to D^{-}T^{\prime\prime+}_{c\bar{s}}$ $\frac{1}{\sqrt{2}}(2(T^{\prime}_{\bar{D}}+P^{\prime}_{\bar{D}})+(C^{\prime}_{\bar{D}T}+P^{\prime}_{T}))$
@@
# Mode $A(\overline{B}_{q}\to\overline{D}T_{cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}})$ # Mode $A(\overline{B}_{q}\to\overline{D}T_{cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}})$
$4^{\prime*}$ $\overline{B}{}^{0}\to D^{-}T^{+}_{cs}$ $\sqrt{2}(C_{\bar{D}T}+P_{T})$ $5^{\prime*}$ $B^{-}\to D^{-}T^{0}_{cs}$ $(C_{\bar{D}T}+P_{T})$
$5^{\prime*}$ $\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{cs}$ $(C_{\bar{D}T}+P_{T})$ $6^{\prime*}$ $B^{-}\to\overline{D}^{0}T^{-}_{cs}$ $\sqrt{2}(C_{\bar{D}T}+P_{T})$
$13^{\prime}$ $\overline{B}{}^{0}\to D_{s}^{-}T^{\prime+}_{c}$ $\frac{\sqrt{3}}{2}(C_{\bar{D}T}+P_{T})$ $13^{\prime}$ $\overline{B}{}_{s}^{0}\to D^{-}T^{\prime+}_{c}$ $\frac{\sqrt{3}}{2}(C_{\bar{D}T}+P_{T})$
$14^{\prime}$ $B^{-}\to D_{s}^{-}T^{\prime 0}_{c}$ $\frac{\sqrt{3}}{2}(C_{\bar{D}T}+P_{T})$ $14^{\prime}$ $\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{\prime 0}_{c}$ $\frac{\sqrt{3}}{2}(C_{\bar{D}T}+P_{T})$
$15^{\prime}$ $\overline{B}{}_{s}^{0}\to D_{s}^{-}T^{\prime+}_{c\bar{s}}$ $(C_{\bar{D}T}+P_{T})$
$16^{\prime}$ $\overline{B}{}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c}$ $\frac{1}{2}(4(T_{\bar{D}}+P_{\bar{D}})+(C_{\bar{D}T}+P_{T}))$ $16^{\prime}$ $\overline{B}{}_{s}^{0}\to D^{-}T^{\prime\prime+}_{c}$ $\frac{1}{2}((C_{\bar{D}T}+P_{T})+4(E_{\bar{D}\bar{B}}+PA_{\bar{B}}))$
$17^{\prime}$ $B^{-}\to D_{s}^{-}T^{\prime\prime 0}_{c}$ $\frac{1}{2}(4(T_{\bar{D}}+P_{\bar{D}})+(C_{\bar{D}T}+P_{T}))$ $17^{\prime}$ $\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{\prime\prime 0}_{c}$ $\frac{1}{2}((C_{\bar{D}T}+P_{T})+4(E_{\bar{D}\bar{B}}+PA_{\bar{B}}))$
$18^{\prime}$ $\overline{B}{}_{s}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c\bar{s}}$ $2(T_{\bar{D}}+P_{\bar{D}})+(C_{\bar{D}T}+P_{T})$
$2(E_{\bar{D}\bar{B}}+PA_{\bar{B}})$
# Mode $A^{\prime}(\overline{B}_{q}\to\overline{D}T_{cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}})$ # Mode $A^{\prime}(\overline{B}_{q}\to\overline{D}T_{cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}})$
$2^{\prime*}$ $\overline{B}{}^{0}\to D_{s}^{-}T^{+}_{c\bar{s}}$ $\frac{(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})}{\sqrt{2}}$ $2^{\prime*}$ $\overline{B}{}_{s}^{0}\to D^{-}T^{+}_{c\bar{s}}$ $\frac{(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})}{\sqrt{2}}$
$3^{\prime*}$ $B^{-}\to D_{s}^{-}T^{0}_{c\bar{s}}$ $(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$ $3^{\prime*}$ $\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{0}_{c\bar{s}}$ $(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$
$8^{\prime*}$ $\overline{B}{}_{s}^{0}\to D_{s}^{-}T^{+}_{c\bar{s}\bar{s}}$ $\sqrt{2}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$
$10^{\prime*}$ $\overline{B}{}^{0}\to D^{-}T^{+}_{c}$ $\sqrt{\frac{2}{3}}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$ $11^{\prime*}$ $B^{-}\to D^{-}T^{0}_{c}$ $\sqrt{\frac{2}{3}}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$
$11^{\prime*}$ $\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{c}$ $\sqrt{\frac{2}{3}}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$ $12^{\prime*}$ $B^{-}\to\overline{D}^{0}T^{-}_{c}$ $\sqrt{2}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$
$13^{\prime}$ $\overline{B}{}^{0}\to D^{-}T^{\prime+}_{c}$ $\frac{1}{\sqrt{3}}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$ $14^{\prime}$ $B^{-}\to D^{-}T^{\prime 0}_{c}$ $\frac{1}{2\sqrt{3}}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$
$14^{\prime}$ $\overline{B}{}^{0}\to\overline{D}^{0}T^{\prime 0}_{c}$ $\frac{1}{2\sqrt{3}}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$
$15^{\prime}$ $\overline{B}{}^{0}\to D_{s}^{-}T^{\prime+}_{c\bar{s}}$ $\frac{1}{2}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$ $15^{\prime}$ $\overline{B}{}_{s}^{0}\to D^{-}T^{+}_{c\bar{s}}$ $\frac{1}{2}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$
$16^{\prime}$ $\overline{B}{}^{0}\to D^{-}T^{\prime\prime+}_{c}$ $2(T^{\prime}_{\bar{D}}+P^{\prime}_{\bar{D}})+(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$ $17^{\prime}$ $B^{-}\to D^{-}T^{\prime\prime 0}_{c}$ $\frac{1}{2}(4(T^{\prime}_{\bar{D}}+P^{\prime}_{\bar{D}})+(C^{\prime}_{\bar{D}T}+P^{\prime}_{T}))$
$2(E^{\prime}_{\bar{D}\bar{B}}+PA^{\prime}_{\bar{B}})$ $17^{\prime}$ $\overline{B}{}^{0}\to\overline{D}{}^{0}T^{\prime\prime 0}_{c}$ $\frac{1}{2}((C^{\prime}_{\bar{D}T}+P^{\prime}_{T})+4(E^{\prime}_{\bar{D}\bar{B}}+PA^{\prime}_{\bar{B}}))$
$18^{\prime}$ $\overline{B}{}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c\bar{s}}$ $\frac{1}{2}(C^{\prime}_{\bar{D}T}+P^{\prime}_{T}+4(E^{\prime}_{\bar{D}\bar{B}}+PA^{\prime}_{\bar{B}}))$ $18^{\prime}$ $\overline{B}{}_{s}^{0}\to D^{-}T^{\prime\prime+}_{c\bar{s}}$ $\frac{1}{2}(4(T^{\prime}_{\bar{D}}+P^{\prime}_{\bar{D}})+(C^{\prime}_{\bar{D}T}+P^{\prime}_{T}))$
@@
From Table 9 we see that the decay modes involving $T$ in $\bf 6$} in scenario I are governed by $(C_{\bar{D}T}+P_{T})$ or $(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$. Consequently, their decay rates are highly related. In particular, the modes involving flavor exotic states have the following relations on decay rates,
$\Gamma_{5}$ $\equiv$ $\Gamma(B^{-}\to D^{-}T^{0}_{cs})=\Gamma(\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{cs}),$
[…]
$\Gamma_{5}$ $\simeq$ $\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma_{6},$ (69)